2. The equation of two regressions obtained in a correlation




analysis of 60 observations are 5𝑥 = 6𝑦 + 24 and




1000𝑦 = 768𝑥 − 3608. What is the correlation coefficient?




Find the ratio of the coefficient of variability of x to that of y.




What is the ratio of the variations of x and y?

1
Expert's answer
2021-11-22T07:21:13-0500

Regression lines:

"x=6y\/5+24\/5"

"y=768x\/1000-3608\/1000"


correlation coefficient:

"r=\sqrt{a_{xy}a_{yx}}=\sqrt{\frac{6}{5}\cdot \frac{768}{1000}}=0.92"


 ratio of the coefficient of variability of x to that of y:


"\frac{V(x)}{V(y)}=\frac{a_{xy}}{a_{yx}}=\frac{768/1000}{6/5}=0.64"


ratio of the variations of x and y:

"r^2=0.92^2=0.8464"